1.Design the longitudinal and shear reinforcement for a 300mm x 400mm column support.The characteristic strength of concrete is 30KN/mm2 and of the steel reinforcement is 460 N/mm2 and 250N/mm2 for the links.The fire resistance requirements is 1 hour for all column.The self weight of concrete is 24KN/M3,the ultimate axial load N=950KN.
Answer #1

Solution:

Given column Dimension: (300x400) mm. fc=30N/mm². fy=460 N/mm² • Fire resistance = 1hr....cie. Nominal cover=40mm). UltmateO . 6 Links !. Provide test i) Diameter of Tie bars 6mm. .- (minimum). • 1/4x0 = V4X22 = 5.smm. Ti) modemum te spacing :). .4